OpenStudy (anonymous):

Complete the square.... y= -4x^2+16x+32

OpenStudy (anonymous):

\[ y= -4 \left( x^2 -4x-8 \right)\\ =-4 \left( x^2 -4x + 4 -4-8 \right)\\ =-4 \left( (x-2)^2 -12 \right)\\ \]